how do i solve k(k-3)=28
Answer by issacodegard(60)

Distribute the k:
k(k-3)=28
k*k-k*3=28
k^2-3k=28
Subtract 28:
k^2-3k-28=0
Factor:
(k-7)(k+4)=0
Since the product of two expressions is zero for some k, one of them must be zero. This implies:
k-7=0 OR k+4=0
This implies:
k=7 OR k=-4